OpenStudy (anonymous):

-12/y=y-7 help please!!

OpenStudy (ankit042):

-12 = y^2-7y solve for y now

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

-12=(y-7)*y y²-7y+12=0 solve for y

OpenStudy (anonymous):

Thank you!!!!!!!